[Hallc_running] [New Logentry] Hall C RC update: 10/19/2021

gaskelld at jlab.org gaskelld at jlab.org
Tue Oct 19 09:40:01 EDT 2021

Logentry Text:
Yesterday's pass change had a relatively modest impact on Hall C with beam gone for about 4 hours (9 am-1 pm).  During the pass change, we took an opportunistic Controlled Access to allow some trouble-shooting of the SHMS HB LHe issues (https://logbooks.jlab.org/entry/3927603).  Also, checks were performed on the new ozone monitor that was installed earlier this month - the readings on the new monitor were shown to agree with a handheld ozone monitor (https://logbooks.jlab.org/entry/3927547).

Beam on late swing and owl shift was fairly unstable due to instabilities in several cryomodules, contributing to relatively low data-taking efficiency for the day overall. For the last 24 hours, we have accumulated an additional ~200 good e,pi-n events.  The last few hours beam has been significantly more stable, averaging about 3-4 trips/hour.

The plan for today is production on LH2 at 42 uA.  Depending on the efficiency over day shift, we may take some data on the 10 cm dummy target on swing shift.

Reminder: There is no RC meeting today.  We will meet again tomorrow at 10 am.


This is a plain text email for clients that cannot display HTML.  The full logentry can be found online at https://logbooks.jlab.org/entry/3928024
-------------- next part --------------
An HTML attachment was scrubbed...
URL: <https://mailman.jlab.org/pipermail/hallc_running/attachments/20211019/f01307d5/attachment.html>

More information about the Hallc_running mailing list